# Finding Visual Task Vectors

### v1

@@ Abstract
Visual Prompting is a technique for teaching models to perform a visual task via in-context examples, without any additional training. In this work, we analyze the activations of MAE-VQGAN, a recent Visual Prompting model [4], and find task vectors, activations that encode task-specific information. Equipped with this insight, we demonstrate that it is possible to identify the task vectors and use them to guide the network towards performing different tasks without providing any input-output examples. To find task vectors, we compute the average intermediate activations per task and use the REINFORCE [43] algorithm to search for the subset of task vectors. The resulting task vectors guide the model towards performing a task better than the original model without the need for input-output examples. [^1] For code and models see www.github.com/alhojel/visual_task_vectors
Figure 1: Visual Prompting models like MAE-VQGAN [4] require an input-output example to describe the desired task in their forward pass. We analyze the model activations and find task vectors, activations that encode task information that can be reused to control the task the model performs (see Figure a). Specifically, we tap into activations of individual attention heads and replace their outputs with task vectors to guide the model to the desired task (see Figure b). Surprisingly, the resulting models perform better than the original model while reducing the need for input-output examples. This confirms that task vectors exist in the network activation space and they can guide the model to perform the desired task.
@@ 1 Introduction
In-context learning (ICL) is an emergent capability of large neural networks, first discovered in GPT-3 [6]. ICL allows models to adapt to novel downstream tasks specified in the user’s prompt. In computer vision, Visual ICL (known as Visual Prompting [4, 2, 3, 51]) is still at its infancy but it is increasingly becoming more popular due to the appeal of using a single model to perform various downstream tasks without specific finetuning or change in the model weights.
@@
In this work, we ask how in-context learning works in computer vision. While this question has yet to be explored, there has been a significant body of research in Natural Language Processing (NLP) trying to explain this phenomenon [1, 7, 13, 14, 15]. Most recently, Hendel et al.[17] suggested that LLMs encode Task Vectors, these are vectors that can be patched into the network activation space and replace the ICL examples while resulting in a similar functionality. Concurrently, Todd et al. [38] discovered Function Vectors, activations of transformer attention heads that carry task representations. Our work is inspired by these observations and aims to study how ICL works in computer vision.
@@
Can task vectors exist in computer vision models too? To build intuition, we start by exploring the activation space of MAE-VQGAN [4]. Intuitively, we look for intermediate activations that are invariant to changes within a task, but have high variance across different tasks. We use this simple metric to rank activations according to their “taskness”. Estimating this score only requires applying the forward pass of a model over minibatches of data and thus it is very fast to compute. Using this score, we find that the activation space of certain attention heads can perfectly cluster the data by tasks, hinting the existence of visual task vectors.
@@
Based on this insight, we hypothesize that Visual ICL models create task vectors too, and aim to find them. However, finding visual task vectors by relying on previous approaches is challenging. For example, both [17, 38] restricted their search space to the output activations of the last token in the prompt sequence (colored in red in the following example: ‘‘Banana:B, Apple:’’ ). This is natural for text as it is processed sequentially by autoregressive models like LLaMA [39], but with images (see Figure 1, “Visual Prompting” input image), it is not obvious what token activations hold task information, and architectures like MAE-VQGAN [4] do not process image tokens sequentially. This alone increases the search space significantly because multiple tokens might hold task vectors.
To find the set of task vectors, we condition the model on query only (see Figure 1a, bottom row) and search for a subset of mean task activations that can be patched in self attention heads to guide the model towards performing the desired task using REINFORCE [43]. Patching the identified task vectors (see example in Figure 1b) leads to competitive performance with the original model on a range of tasks, which confirms that the set of resulting activations are indeed task vectors. Our approach directly models searching for a subset of activations compared to the previously proposed Casual Mediation Analysis [32] (CMA) that searches each activation independently.
@@
Our contributions are as follows. We show evidence for the existence of visual task vectors and propose a practical way to identify them. Moreover, we find that patching the resulting task vectors guides the model towards the desired task with similar performance compared to the original model while reducing the number of FLOPS by $22.5\%$.
@@ 2.2 Explainability
In the context of enhancing model interpretability [50, 29, 37] within computer vision, the integration of Causal Interventions [5, 31] and Activation Patching [48] has become a cornerstone in the elucidation of complex neural networks’ decision-making processes. These methodologies enable a systematic examination of how models encode and utilize high-level concepts, offering profound insights into their internal mechanisms[49, 44, 24]. Causal Interventions [32, 28], facilitate the exploration of the causal structures of models by manipulating their internal states or inputs [12] and observing the impact on outputs, thus uncovering the direct and indirect effects that drive model predictions. Here we use Activation Patching [48] to show through targeted interventions the significance of Task Vectors towards guiding Visual Prompting models to perform different computer vision tasks.
@@ 2.3 Task Vectors
In [17, 38, 11], a Task Vector or Function Vector is a latent activation derived from a particular layer of the transformer [40]. This vector subsequently substitutes the original latent states at the same layer during the forward pass of a query to guide the model to perform the desired task. The investigation of Task Vectors aligns with broader efforts in the field to make neural networks more adaptable and tailored to specific tasks [23, 26] as well as boosting the performance [30, 47, 19] by gaining a deeper understanding of how different layers within a model contribute to its overall function. Our work is the first to explore task vectors in computer vision.
@@ 3 Methods
Our goal is to understand in-context learning for computer vision, and how existing models can be adapted to different downstream task in inference time. Specifically, we focus on the MAE-VQGAN [4] model, a variant of MAE [16] with a Vision Transformer [9] encoder-decoder architecture.
Given a an input-output example $(x_{s},y_{s})$ and a new query $x_{q}$, to succeed in this ICL task, the model $F$ has to implicitly apply the transformation from $x_{s}$ to $y_{s}$ over $x_{q}$ to produce $y_{q}$:
@@
Based on past observations from NLP [17, 38], we hypothesize that computer vision models encode latent task vectors in their activation space during the forward pass as well. This requires the model to implicitly map the input example into a set of latent task vectors $z=\{z_{i}\}^{k}_{i=1}$. Then, the original function $F$ can be decomposed into extracting the task vectors by a function $G$ then applying $F$ on the query while fixing the computed task activations $z$:
@@ 3.1 Scoring Activations
Intuitively, every task vector is an activation that changes across different tasks but remains invariant to changes within a specific task. Specifically, denote $i=(l,m,k)$ as the $l^{th}$ attention block, $m^{th}$ attention head, and $k^{th}$ token and denote it by $F_{i}$ the function that outputs the corresponding activation in the attention block residual stream. We sample an input-output example and a query triplet from individual tasks as well as across tasks and compute the latent activations corresponding to $i$:
$(x_{s},y_{s},x_{q})\sim\mathbf{D_{all\_tasks}}$ $(x^{\prime}_{s},y^{\prime}_{s},x^{\prime}_{q})\sim\mathbf{D_{task\_j}}$
$z^{i}_{all}=F_{i}(x_{s},y_{s},x_{q})$ $z^{i}_{task_{j}}=F_{i}(x^{\prime}_{s},y^{\prime}_{s},x^{\prime}_{q})$
Then, we define the score and the mean activations for position $i$ and task $j$:
${\rho_{token}}(i)=\frac{var(z^{i}_{all})}{{\frac{1}{n}}\sum^{n}_{j=1}var(z^{i}_{task_{j}})}$ $\mu_{i,j}=\mathbb{E}[z^{i}_{task_{j}}]$ (3)
@@
Table 1: Task Clustering Quality. Clustering Scores of Different Attention Heads, ranked by our Activation Scoring metric (see Section 3). This indicates that higher Activation Scores indeed correlate with better clustering by tasks.

### v2

@@ Abstract
Visual Prompting is a technique for teaching models to perform a visual task via in-context examples, without any additional training. In this work, we analyze the activations of MAE-VQGAN, a recent Visual Prompting model [4], and find Task Vectors, activations that encode task-specific information. Equipped with this insight, we demonstrate that it is possible to identify the Task Vectors and use them to guide the network towards performing different tasks without having to provide any in-context input-output examples. To find Task Vectors, we compute the mean activations of the attention heads in the model per task and use the REINFORCE [43] algorithm to patch into a subset of them with a new query image. The resulting Task Vectors guide the model towards performing the task better than the original model. [^1] For code and models see www.github.com/alhojel/visual_task_vectors
Figure 1: Visual Prompting models like MAE-VQGAN [4] require input-output example(s) to describe the desired task in their forward pass. We analyze the model activations and find Task Vectors, activations that encode task information that can be reused to control the task the model performs (see Figure 1a). Specifically, we tap into activations of individual attention heads and replace their outputs with Task Vectors to guide the model to the desired task (see Figure 1b). Surprisingly, the resulting models perform better than the original model while removing the need for input-output examples. This confirms that Task Vectors exist in the network activation space and they can guide the model to perform the desired task.
@@ 1 Introduction
In-context learning (ICL) is an emergent capability of large neural networks, first discovered in GPT-3 [6], which allows models to adapt to novel downstream tasks specified in the user’s prompt. In computer vision, Visual ICL (known as Visual Prompting [4, 2, 3, 51]) is still at its infancy but it is increasingly becoming more popular due to the appeal of using a single model to perform various downstream tasks without specific finetuning or change in the model weights.
@@
In this work, we ask how in-context learning works in computer vision. While this question has yet to be explored, there has been a significant body of research in Natural Language Processing (NLP) trying to explain this phenomenon [1, 7, 13, 14, 15]. Most recently, Hendel et al.[17] suggested that LLMs encode Task Vectors, these are vectors that can be patched into the network activation space and replace the ICL examples while resulting in a similar functionality. Concurrently, Todd et al. [38] discovered Function Vectors, activations of transformer attention heads that carry task representations. Our work is inspired by these observations and we aim to study ICL in computer vision.
@@
Do Task Vectors exist in computer vision models as well? To build intuition, we start by exploring the activation space of MAE-VQGAN [4]. Intuitively, we look for intermediate activations that are invariant to changes within a task, but have high variance across different tasks. We use this simple and fast-to-compute metric to rank activations according to their “taskness”. Using this score, we find that the activation space of certain attention heads can perfectly cluster the data by tasks, hinting at the existence of Visual Task Vectors.
@@
Based on this insight, we hypothesize that Visual ICL models create Task Vectors too, and aim to find them. However, finding Visual Task Vectors by relying on previous approaches is challenging. For example, both [17, 38] restricted their search space to the output activations of the last token in the prompt sequence (colored in red in the following example: ‘‘Banana:B, Apple:’’ ). This approach is natural for text as it is processed sequentially by autoregressive models like LLaMA [39], but with images (see Figure 1a, “Visual Prompting” input image), it is not obvious what token activations hold task information, and architectures like MAE-VQGAN [4] do not process image tokens sequentially. This alone increases the search space significantly because multiple tokens might hold Task Vectors. Additionally, while Hendel et al. [17] and Todd et al. [38] differ in the activation space in which they perform interventions, both methods evaluate single interventions, ignoring emergent second-order effects.
To identify Task Vectors, we first compute the mean attention head outputs for each token position across a set of task examples (Figure 1a). We then replace attention head outputs at selected positions with these pre-calculated means, using a new query as input. Each position is modeled as a random variable, and we use REINFORCE [43] to optimize patching positions, guiding the model towards the desired task. This approach yields competitive performance across various tasks, validating the existence of Task Vectors. Our method effectively explores higher-order effects and complex interactions between attention heads in different layers, improving upon previous approaches that examine activations in isolation.
@@
Our contributions are as follows. We show evidence for the existence of Visual Task Vectors and propose a practical way to identify them. Moreover, we find that patching the resulting Task Vectors guides the model towards the desired task with better performance compared to the original in-context prompts while reducing the number of FLOPS needed for a forward pass by $22.5\%$.
@@ 2.2 Explainability
Causal Interventions [5, 31, 32, 28] and Activation Patching [48] are valuable tools for understanding complex neural networks’ internal mechanisms, enhancing model interpretability [50, 29, 37]. These methods enable systematic examination of how models encode information and represent high-level concepts [49, 44, 24]. By manipulating internal states or inputs and observing output changes, they reveal causal structures and effects driving model predictions [12]. In this work, we employ Activation Patching [48] to improve Visual Prompting models’ guidance for various computer vision tasks through targeted interventions.
@@ 2.3 Task Vectors
In [17, 38, 11], Task Vectors or Function Vectors are sets of latent activations derived from particular positions inside a transformer [40] model which serve as internal representations of a task implicitly described by an ICL prompt with input-output demonstrations. These latent activations at particular positions can consequently be used in a new forward pass in the absence of the ICL prompt (or with a corrupted prompt) while still managing to guide the model to perform the desired task. The investigation of Task Vectors aligns with broader efforts in the field to make neural networks more adaptable and tailored to specific tasks [23, 26] as well as boosting the performance [30, 47, 19] by gaining a deeper understanding of how different attention heads within a model contribute to its overall function. Our work is the first to explore Task Vectors in computer vision.
@@ 3 Methods
Our goal is to understand in-context learning for computer vision and how existing models can be adapted to different downstream tasks at inference time. We focus on the MAE-VQGAN model [4], a variant of MAE [16] with a Vision Transformer [9] encoder-decoder architecture. Given an input-output example $(x_{s},y_{s})$ and a new query $x_{q}$, to succeed in an ICL task, the model $F$ must implicitly apply the transformation from $x_{s}$ to $y_{s}$ over $x_{q}$ to produce $y_{q}$:
@@
Based on observations from NLP [17, 38], we hypothesize that computer vision models also encode latent Task Vectors in their activation space during the forward pass. This requires the model to implicitly map the ICL example into a set of latent Task Vectors $z=\{z_{i}\}^{k}_{i=1}$ which we derive from the attention head outputs for different tokens across the model’s layers (the internal representations). The original function $F$ can then be decomposed into extracting the Task Vectors by a function $G$ and applying $F$ on the query while fixing the computed task activations $z\in\mathbb{R}^{d}$, where $d$ is the hidden dimension of the model.
@@
We now describe how Task Vectors can be derived from the internal representations of a model.
3.1 Computing Mean Activations
Let $i=(l,m,k)$ denote the position in the model, where $l$, $m$, and $k$ are the attention block, head, and token indices respectively. Define $H_{i}:(x_{s},y_{s},x_{q})\rightarrow\mathbb{R}^{d}$ as the function that outputs the activation at position $i$ for a given demonstration and query triplet. Let $(x_{s},y_{s},x_{q})\sim\mathbf{D_{task_{j}}}$ be a triplet of input-output example and a query from task $j$. We compute the intermediate activation $h^{i}_{task_{j}}$ via $H_{i}$ and denote its mean activation as $\mu_{i,j}$:
$h^{i}_{task_{j}}=H_{i}(x_{s},y_{s},x_{q})$ $\mu_{i,j}=\mathbb{E}[h^{i}_{task_{j}}]$ (3)
@@ 3.2 Scoring Activations
Intuitively, every task vector is an activation that changes across different tasks but remains relatively invariant to changes within a specific task. We define the data distribution $(x^{\prime}_{s},y^{\prime}_{s},x^{\prime}_{q})\sim\mathbf{D_{all\_tasks}}$ as the union of all task-specific distributions, and the intermediate activations $h^{i}_{all}=H_{i}(x^{\prime}_{s},y^{\prime}_{s},x^{\prime}_{q})$. The scoring function $\rho_{token}(i)$ is defined as the ratio of inter-task to intra-task variance as $\rho_{token}(i)$ where $\text{Var}(\cdot)$ denotes the variance, $h[e]$ is the $e$-th element of vector $h$, and $n$ is the number of tasks.
$\rho_{token}(i)$ $=\frac{\sum_{e=1}^{d}\text{Var}(h^{i}_{all}[e])}{\frac{1}{n}\sum_{j=1}^{n}\sum_{e=1}^{d}\text{Var}(h^{i}_{task_{j}}[e])}$ (4)
Computing these is efficient, requiring only forward passes across batches of data. This scoring function identifies "taskness" in activations (see Figure 2, and Table 1), but may also highlight other task-variant activations (e.g., color histograms varying across tasks but consistent within tasks like binary segmentation maps versus colored images). We use these scores mainly to build intuition, analyze task clusters in the activation space (Section 5), and develop a simple baseline (“Greedy Random Search”, see Suppl. 11.1). However, we employ a more robust general approach to identifying high-order Visual Task Vectors, as described in the following section.
3.3 Finding Visual Task Vectors via REINFORCE
For every task $j$ we can apply the following procedure to find the Task Vectors. Recall that $\{\mu_{i,j}\}$ denotes the mean activations and $F(\cdot)$ is a pretrained visual prompting model. Denote $\alpha_{i,j}\sim Bernoulli(\sigma(\theta_{ij}))$ as random variables that signify whether the mean task activation $\mu_{i,j}$ is a task vector of task $j$ placed in activation position $i$ and $\theta_{ij}$ is a learned weight followed by the sigmoid function to ensure it is in $[0,1]$.
Denote $z_{j}$ as the set of Task Vectors for task $j$: $z_{j}=\{\mu_{i,j}|\alpha_{i,j}=1\}$. Given pairs of input-output task demonstrations $x_{q},y_{q}\sim\mathbf{D_{task_{j}}}$, [^2] We overload the definition of $\mathbf{D_{task_{j}}}$ to avoid notation clutter. we want to find the set of Task Vectors that minimizes the loss function:
$L(\theta)=E_{{z_{j}}\sim p_{\theta}}L(y_{q},F(x_{q}|z_{j}))$ (5)
Figure 2: Activation Scoring Analysis. Individual scores (${\rho_{token}}(i)$) aggregated per Attention Head (left) for the encoder and decoder. Individual token scores of specific heads, along with the t-SNE [27] clustering of head activations of different tasks (right).
Instead of learning Task Vectors placements ${\alpha_{ij}}$ for each individual task, we can revise the procedure above to learn a single placement by defining random variables $\alpha_{i}\sim Bernoulli(\sigma(\theta_{i}))$ that accommodate all tasks. In this setting, training requires drawing examples across tasks with their respective task-wise mean activations and task specific loss function. We refer to this setting as multi-task patching. Furthermore, the search can be applied at different granularities, such as patching token groups from the same quadrant, all tokens in an attention head, or entire layers. We discuss these design choices in the next section.
4 Experiments
Our experiments explore if activation patching of Task Vectors can make a model perform the desired visual task as well as or better than its original one-shot performance. We describe the implementation of MAE-VQGAN, prompting schemes, baselines for comparison, visual tasks, and conducted experiments.
4.1 Implementation Details
MAE-VQGAN [4]. An MAE  [16] with a ViT-L [9] backbone. The decoder predicts a distribution over a VQGAN [10] codebook to output images with better visual quality. We used the pretrained checkpoint from [4] where it was trained over the Computer Vision Figures [4] dataset and ImageNet [35].
Finding Task Vectors. Similar to one-shot (Section 4.4), we use a 2x2 image grid with the prompt. For task vectors, we embed only the query in the bottom left quadrant. The model reconstructs only the output part (bottom right). We patchify the query image at 112x112 resolution, apply bottom left quadrant positional encodings, feed patches into the encoder, and process them with the decoder alongside bottom right quadrant mask tokens to obtain the result.
We intervene on attention head outputs by replacing them with mean activations at positions identified using REINFORCE [43]. Initially, we set $\theta_{ij}=-1$. To reduce the search space, we group patching positions into three categories: CLS (1 token), bottom left quadrant (49 query image patch tokens), and bottom right quadrant (49 MASK tokens). In each iteration, we sample 32 times from the Bernoulli distribution for each of 10 images, patching positions where the sampled value is 1. This results in 320 executions of task vector conditioned MAE-VQGAN per iteration. We optimize Bernoulli parameters as outlined in Section 3.3 using Adam [20] with a learning rate of 0.1. The algorithm runs for 600 steps, selecting the best checkpoint every 50 steps based on evaluation on a held-out test set.
4.2 Activation Scoring Analysis
Here our goal is to evaluate the Activation Scoring step (outlined in Section 3.2), specifically whether high scoring activations indeed correspond to Task Vectors.
Collecting Activations. To compute activation scores, we first run the model’s forward pass in a one-shot setting across different tasks (Section 3.1). We use 100 prompts and queries from Pascal 5i [36] training set, ensuring reasonable one-shot performance. We save the activations for each task $j$ and position $i=(l,m,k)$, where $l$, $m$, and $k$ are the attention block, head, and token indices respectively. We then compute the mean activation ${\mu_{i,j}}$ and score $\rho_{token}(i)$.
4.3 Downstream Tasks
Dataset. We utilize Pascal-5i [36], consisting of 4 image splits (346-725 images each) with segmentation masks. For evaluation, We sample 1000 prompt-query pairs per split from the validation set. For Activation Scoring and methods to find Task Vectors, we use only training set examples.
Foreground Segmentation. We use Pascal-5i [36] segmentation masks and report mean IOU (mIOU) across four splits.
Low Light Enhancement. We multiply Pascal-5i image color channels by 0.5 for input, using the original as output. We report Mean Squared Error (MSE).
Inpainting. We mask a random 25% square region (1/8 area) of each image for input. We report MSE using the original image as output.
4.4 Baselines
4.5 Ablations
In this section, we describe the set of experiments conducted to validate our implementation choices of our REINFORCE [43] method.
Task Vectors Location in Encoder vs. Decoder. We hypothesize that task implementation spans both encoder and decoder. To test this, we apply interventions to the encoder only, decoder only, and the whole network. We report mIoU on four Segmentation task splits to assess the necessity of interventions in both parts for effective task implementation.
Patching Granularity. We explore intervention granularities by grouping token positions to reduce dimensionality and optimization search space. Positions $i=(l,m,k)$ are grouped into spatial quadrants or attention heads. To balance precision and search space, we use three granularity levels: individual tokens, quadrants, and attention heads, and report performance across four tasks.
4.6 Task Vector Patching
We investigate whether patching Task Vectors into a model’s forward pass, without ICL demonstrations, can achieve performance comparable to one-shot prompting. Our experiments include:
Multi-task. We apply our method in a multi-task scenario, selecting two images per task and evaluating alongside CMA and GRS baselines. We jointly assess loss across all tasks, normalizing per task to ensure equal weighting and prevent single-task dominance. We include an identity-copy task to maintain a consistent batch size of 10 for fair comparisons.
MAE-VQGAN. We compare to the original model’s one-shot performance.
GRS across Random K Layers. We execute the search algorithm on a random set of $k$ layers, instead of score-ranked layers, to validate the necessity of our proposed scoring method.
We report results across four splits for all four tasks, comparing our variants to the original MAE-VQGAN model and random baselines. Supplementary Materials (Section 4) include initial explorations of REINFORCE-powered Language Task Vectors with Llama7B, vector addition to create composable tasks, and others.
5 Results
5.1 Activation Scoring Analysis
We compute and display ${\rho_{token}}(i)$ per head on a heatmap, with layers on the y-axis and head indices on the x-axis (Figure 2). This highlights heads that potentially hold Task Vectors. We select top heads—$(26,3)$ and $(11,3)$—and a lower-ranked head $(27,5)$. For each, we visualize activation clustering and individual Activation Score per token.
Clustering Visualization. Highly-ranked heads (e.g., (26,3)) show clear task-based clustering, while low-ranked heads like (27,5) show many small clusters with different tasks, likely based on input semantics. The activations are projected onto 2D using t-SNE [27] with colors indicating tasks.
Score-per-token Heatmap. We display ${\rho_{token}}(i)$ values for each token, reflecting spatial positioning in a 2x2 grid. For heads of interest, these values are shown on a heatmap. The CLS token is in the top left, followed by values for quadrants $(x_{s},y_{s},x_{q},y_{q})$. The encoder lacks $y_{q}$ tokens (bottom right), marked as X. Token values within a head vary widely but show consistency within quadrants, supporting quadrant-based token patching.
Quantitative Clustering Analysis. To validate our scoring-based clustering quality observations, we report Silhouette and Davies-Bouldin scores for the two highest-ranked heads, two randomly sampled heads, and the two lowest-ranked heads (Table 1). Heads scored highly by our method show high-quality clustering scores, while low-scored heads show poor clustering. Randomly selected heads received intermediate scores, further supporting our methodology.
@@ 5.1 Activation Scoring Analysis
Table 1: Task Clustering Quality. Clustering Scores of Different Attention Heads, ranked by our Activation Score (see Section 3). This indicates that higher Activation Scores indeed correlate with better clustering by tasks.